# Towards the New XAI: A Hypothesis-Driven Approach to Decision Support Using Evidence

### v3

@@ 6 Discussions
Strengths and weaknesses of our hypothesis-driven approach
First, participants using the hypothesis-driven approach required a similar time to complete the task compared to the recommendation-driven approach. Participants in the hypothesis-driven condition also made higher quality decisions than recommendation-driven and AI explanation only based on the Brier score. The results indicated that the hypothesis-driven gave study participants a more complete picture of the underlying decision aid than the other two approaches, helping them to make use of the AI models when they are right, and be less confident when the models are wrong.
Moreover, hypothesis-driven reduced over-reliance significantly compared to the standard AI recommendation. Similarly, hypothesis-driven also reduced under-reliance compared to AI explanation only. Importantly, the positive result for under-reliance using recommendation-driven is not cancelled out by the poor over-reliance result, compared to hypothesis-driven. The primary aim indicating potential for the use of uncertainty/confidence [5, 24] and conformal prediction [33] to direct decision makers’ attention towards a set of hypotheses that it is confident about.
@@
On the other hand, recommendation-driven and AI explanation only do not support this. We found that in recommendation-driven, people could use feature values to confirm the validity of the decision aid’s prediction. However, they are not aware of the uncertainty among different hypotheses. In AI explanation only, people often ignore using the evidence and solely focus on using the feature values to make a decision because interpreting the evidence with this approach can be a lot more mentally demanding.
Study limitations
@@
There are also some limitations with the study. First, we ran the experiment on one dataset (Ames Housing), which limits generalisability. In addition, as there is no ground truth for the price of a house, the experimental participants’ tasks are somewhat subjective. Further, this task has only three output classes, so only three hypotheses, and we anticipate the results would be more interesting when we consider more hypotheses. Finally, the human experiment is currently conducted with laypeople while experts likely interact with the decision-aiding tool differently from laypeople [14].
7 Conclusions
In this paper, we show that the hypothesis-driven approach using Weight of Evidence (WoE) can significantly reduce reliance, improve decision-making quality compared to two other prevalent decision-making approaches (recommendation-driven and AI explanation only). Furthermore, hypothesis-driven helps participants to be aware of the uncertainty among multiple options. Nevertheless, there still remains a challenge of study participants relying on the wrong (or misleading) evidence. Therefore, future work can address this challenge by exploring different approaches for presenting trustworthy evidence. More generally, potential future work is to consider the uncertainty/confidence in the generated evidence.
{ack}
This research was supported by the University of Melbourne Research Scholarship (MRS) and partly funded by Australian Research Council Discovery Grant DP190103414. We also want to thank the anonymous reviewers for their valuable feedback.
Appendix A Statistics of Experiment 1
count mean std min 25% 50% 75% max
condition (median)
(C1) Recommendation-driven 102.000 17.920 9.342 5.100 11.492 15.767 21.792 67.017
(C2) AI-explanation-only 99.000 18.619 11.258 3.500 12.375 15.467 22.417 92.650
(C3) Hypothesis-driven 101.000 18.087 9.255 5.967 11.850 15.733 21.450 55.833
Table 2: Statistics of completion time per condition (in minutes).
count mean std min 25% 50% 75% max
condition (median)
(C1) Recommendation-driven 102.000 0.290 0.071 0.183 0.239 0.277 0.333 0.484
(C2) AI-explanation-only 99.000 0.295 0.073 0.140 0.242 0.281 0.337 0.594
(C3) Hypothesis-driven 101.000 0.267 0.063 0.154 0.228 0.252 0.304 0.474
Table 3: Statistics of Brier score per condition.
count mean std min 25% 50% 75% max
condition (median)
(C1) Recommendation-driven 102.000 73.856 20.913 33.333 66.667 66.667 100.000 100.000
(C2) AI-explanation-only 99.000 54.209 22.505 0.000 41.667 50.000 66.667 100.000
(C3) Hypothesis-driven 101.000 53.300 22.733 16.667 33.333 66.667 66.667 100.000
Table 4: Statistics of over-reliance (%) per condition.
count mean std min 25% 50% 75% max
condition (median)
(C1) Recommendation-driven 102.000 17.810 20.346 0.000 0.000 16.667 33.333 100.000
(C2) AI-explanation-only 99.000 41.246 27.183 0.000 16.667 33.333 50.000 100.000
(C3) Hypothesis-driven 101.000 24.422 18.191 0.000 16.667 16.667 33.333 100.000
Table 5: Statistics of under-reliance (%) per condition.
In this section, we show the statistics of four measures in Experiment 1 as in Table 2,  3,  4 and  5. We also show the boxplot of these measures in Figure 5,  6,  7 and  8.
Figure 5: Completion time. Lower is better. Means represented as dots.
Figure 6: Brier score. Lower is better. Means represented as dots.
Figure 7: Over-reliance. Lower is better. Means represented as dots.
Figure 8: Under-reliance. Lower is better. Means represented as dots.
Appendix B Subjective Questions in Experiment 1
After doing the experiment 1, participants were asked to answer 12 subjective questions as follows.
  1. 1.
    In control: I feel in control of the decision-making process when using this decision aid. (0 = Disagree strongly; 10 = Agree strongly)
  2. 2.
    Preference: I would like to use this decision aid frequently. (0 = Disagree strongly; 10 = Agree strongly)
  3. 3.
    Mental demand: I found this task difficult. (0 = Disagree strongly; 10 = Agree strongly)
  4. 4.
    System complexity: The decision aid was complex. (0 = Disagree strongly; 10 = Agree strongly)
  5. 5.
    Trust: I am confident in the decision aid. I feel that it works well. (0 = Disagree strongly; 10 = Agree strongly)
  6. 6.
    Trust: The decision aid is very predictable. (0 = Disagree strongly; 10 = Agree strongly)
  7. 7.
    Trust: The decision aid is very reliable. I can count on it to be correct all the time. (0 = Disagree strongly; 10 = Agree strongly)
  8. 8.
    Trust: I feel safe that when I rely on the decision aid I will get the right answers. (0 = Disagree strongly; 10 = Agree strongly)
  9. 9.
    Trust: The decision aid is efficient in that it works very quickly. (0 = Disagree strongly; 10 = Agree strongly)
  10. 10.
    Trust: I am wary of the decision aid. (0 = Disagree strongly; 10 = Agree strongly)
  11. 11.
    Trust: The decision aid can perform the task better than a novice human user. (0 = Disagree strongly; 10 = Agree strongly)
  12. 12.
    Trust: I like using the decision aid for decision making. (0 = Disagree strongly; 10 = Agree strongly)
We evaluate Q1-4 separately to measure 4 measures (In control, Preference, Mental demand and System complexity). We aggregate Q5-12 to measure Trust. We based the trust questions on questions from [18].
Based on Figure 9,  10,  11,  12 and  13, AI-explanation-only is significantly worse than the other conditions in all facets. Moreover, recommendation-driven and hypothesis-driven are quite similar and there is no statistical difference between these two conditions. The reason is that we conduct between-subject experiments so each participant has access to only one condition and they do not have another condition to compare to. If we run within-subject experiments to measure subjective questions in the future, participants can compare different decision-making approaches and evaluate which one they prefer the most.
Figure 9: In control (Higher is better)
Figure 10: Preference (Higher is better)
Figure 11: Mental demand (Lower is better)
Figure 12: System complexity (Lower is better)
Figure 13: Trust (Higher is better)
Appendix C Example Questions in Human Experiment
In this section, we will show example questions in our human experiments. There are two main phases in each experiment:
  • Training phase: Participants were given three example questions.
  • Test phase: After finishing the training phase, participants were given twelve test questions.
First, Figure 1416 and 18 show the first page of the training phase in three conditions. This page introduces the task, how to read the evidence, what information will be given and how the participants will be scored. After that, three example questions will be shown, one of which is Figure 15 or 17 or 19 depending on the condition.
Figure 14: Introduction page in (C1) Recommendation-driven
Figure 15: Training phase in (C1) Recommendation-driven: An example question
Figure 16: Introduction page in (C2) AI-explanation-only
Figure 17: Training phase in (C2) AI-explanation-only: An example question
Figure 18: Introduction page in (C3) Hypothesis-driven
Figure 19: Training phase in (C3) Hypothesis-driven: An example question
Second, after finishing the training phase, there are twelve questions in the test phase. Here are the following tasks in this phase.
  • Experiment 1: Assign the likelihood for each price range (Low/Medium/High) of the given house
  • Experiment 2: In addition to the same tasks as in experiment 1, participants were asked to explain their decision using free text. Participants in Experiment 1 were not allowed to participate again in Experiment 2.
Example questions in Experiment 1 are shown in Figure 20, 21 and 22. Experiment 2 will use the same set of questions, with an addition of having a free text after the likelihood slider to ask participants why they made such choices.
To compare between the three conditions (C1) Recommendation-driven, (C2) AI-explanation-only and (C3) Hypothesis-driven:
  • In Figure 20, participants can see the AI prediction (i.e. low price in this case) and the weight of evidence (the explanation) for that prediction.
  • In Figure 21, the AI prediction is hidden. Therefore, even though the participants can see the explanation, they do not know which class (low/medium/high) the evidence refers to.
  • In Figure 22, the house features selected are similar to the example question in Figure 20 and 21. We show participants the evidence for all hypotheses (low, medium and high). We do not give them the AI prediction. Specifically, we have supportive evidence in most features for hypothesis low. By contrast, strongly negative evidence refute hypothesis high. The correct answer here is low.
Figure 20: A screenshot of a question in (C1) Recommendation-driven.
Figure 21: A screenshot of a question in (C2) AI-explanation-only.
Figure 22: Screenshots of evidence provided for all three hypotheses in (C3) Hypothesis-driven.

### v4

@@ 6 Discussions
In this section, we summarise the findings from the experiments and discuss the implications of the results. We also discuss the validity, limitations of the studies and suggest future work.
6.1 The Two Sides of the Coin: Recommendation-driven and Hypothesis-driven
6.1.1 Housing price prediction domain
Participants using the hypothesis-driven approach required a similar time to complete the task compared to the recommendation-driven approach. Participants in the hypothesis-driven condition also made higher quality decisions than recommendation-driven and AI-explanation-only based on the Brier score. The results indicated that the hypothesis-driven gave study participants a more complete picture of the underlying decision aid than the other two approaches, helping them to make use of the AI models when they are right, and be less confident when the models are wrong.
Moreover, the hypothesis-driven approach reduced over-reliance significantly compared to the standard AI recommendation. Similarly, it also reduced under-reliance compared to AI-explanation-only. Importantly, the positive result for under-reliance using the recommendation-driven approach is not cancelled out by the poor over-reliance result, compared to the hypothesis-driven. The primary aim indicates potential for the use of uncertainty/confidence [7, 50] and conformal prediction [75] to direct decision-makers’ attention towards a set of hypotheses that it is confident about.
@@ 6.1.1 Housing price prediction domain
On the other hand, recommendation-driven and AI-explanation-only do not support this. We found that in recommendation-driven, people could use feature values to confirm the validity of the decision aid’s prediction. However, they were not aware of the uncertainty among different hypotheses. In AI-explanation-only, people often ignore using the evidence and solely focus on using the feature values to make a decision because interpreting the evidence with this approach can be a lot more mentally demanding.
6.1.2 Skin cancer diagnosis domain
Recommendation-driven is preferred by beginners who do not have much experience in the field. It is easier to use and faster to make a decision because they only need to confirm if they want to follow the recommendation or not. However, it can bias the clinicians and make them doubt their own diagnosis.
Hypothesis-driven is more suitable for experienced participants who have knowledge in the field. It allows users to compare and contrast between different hypotheses and check where the AI is looking for each diagnosis. However, it requires users to have a good base of knowledge for all possible hypotheses to be able to evaluate the provided evidence. It can also be overwhelming for users who do not have much experience in the domain. In addition, the evidence provided by the AI can be wrong, which can lead to a wrong decision.
Moreover, the order of AI in relation to the human decision-maker is also important. We can either put the AI first or the human first, both of which have their own advantages and disadvantages [31]. The AI-first approach can be used as a triage tool to reduce the assessment time, but it can face regulatory challenges and over-reliance on AI. The human-first approach can retain the current clinical workflow and use AI as a second opinion, which can increase the sensitivity of the diagnosis. However, it can be time-consuming, with a potential increase in consultation time when there are disagreements between the AI and the human. In our study, recommendation-driven implemented the AI first, while hypothesis-driven could be improved by putting the human opinions first before providing the evidence from the AI.
6.2 Laypeople and Skilled Individuals
We conducted two human experiments in two different domains: housing price prediction and skin cancer diagnosis. The goal is to have a more diverse set of participants, including both laypeople and skilled individuals.
In the housing price prediction domain, we found that the hypothesis-driven approach can help participants make better decisions. Note that there were only three possible hypotheses in the first domain, whereas there were seven in the second domain. In case of having too many options, the hypothesis-driven approach can be overwhelming for participants, especially for laypeople or beginners in the field. In the second domain, we found that experienced participants applied their domain knowledge to narrow down possible hypotheses and improve their decision-making overall.
In both studies, participants agree that hypothesis-driven is more informative. However, the recommendation-driven approach is more straightforward and easier to use. In fact, some skilled individuals prefer the recommendation-driven approach because they already have a good base of knowledge to make the decision. They can use the AI to simply compare and contrast their own thought with the AI’s recommendation, leading to a faster decision-making process. In case of having high uncertainty in the decision, hypothesis-driven is preferred by everyone as it provides more information.
6.3 Threats to Validity
We will identify threats to the validity of the human studies, including both internal and external validity.
6.3.1 Housing Price Prediction Domain
@@
There is a threat to internal validity due to instrumentation, we ran the experiment on one dataset (Ames Housing), which limits generalisability. In addition, since we follow the labels from this particular dataset, there is no single ground-truth for the price of a house. The price can vary depending on many factors. Therefore, the experimental participants’ tasks are somewhat subjective. Further, this task has only three output classes, so only three hypotheses, and we anticipate the results would be more interesting when we consider more hypotheses. The current form of explanation can be further improved by incorporating contrastive explanations. At present, we present multiple WoEs for different hypotheses, but these cannot be considered as contrastive explanations.
Regarding the experimental design, there are multiple forms of cognitive forcing, and we tested just one. We could have added another experimental condition where we asked participants to first think about their decision before showing the AI recommendation. This condition may have close performance to the hypothesis-driven approach. Finally, the human experiment is currently conducted with laypeople while experts likely interact with the decision-aiding tool differently from laypeople [21]. Another limitation is that we currently use the output labels of the model as the hypotheses, while domain experts may have more and different hypotheses.
6.3.2 Skin Cancer Diagnosis Domain
Threats to Internal Validity
A threat to internal validity for a long human experiment (i.e., required approximately one hour to finish) is the maturation. Participants could become tired over time and lose concentration when doing later tasks. Instrumentation is the next threat that needs to be considered. We use labels from the HAM10000 dataset [78] to measure the performance of participants. But it is important to note that there is no single ground-truth. Different experts in the field can still have different opinions on the labels. Therefore, we will need a better approach to measuring the correctness of diagnoses rather than solely relying on these labels. Moreover, our pool of participants is relatively small (14 participants), the number of tasks is limited (16 tasks), trained on a single dataset (HAM10000), and limited number of conditions as we did not consider no AI condition in the human experiment. All these factors can affect the significance of the results.
Threats to External Validity
The first threat to external validity is sample characteristics. We focus on using our network to invite experienced diagnosers to participate. This may not generalise to a broader population. Additionally, our participants are not experts. They have background knowledge in the field of skin cancer but are still far from being experts. In our current user study, we categorised them into experienced (those who have received clinical training) and inexperienced (those who have not received clinical training) participants. Future work should consider recruiting experts in the field, which could provide further insights into how domain expertise might influence reliance on AI.
Secondly, regarding ecological validity, laboratory experiments very often do not reflect the real world. For example, while some images can be straightforward, doctors typically require additional information such as the history of the patient, age, sex, regional images, etc. before making a diagnosis. Relying on a single dermoscopic image can be challenging for participants and may lead to high diagnostic uncertainty. Moreover, there is a concern that while AI can perform extremely well in a controlled environment, their behaviour may differ substantially in clinical settings. That is why none of our experienced participants have used AI to support skin cancer diagnosis in clinical practice. At this stage, there is no formally accredited AI system available, and AI is used solely for research and internal testing purposes.
6.4 Future Work
6.4.1 Combining Recommendation-driven and Hypothesis-driven
The question now is how should we design the decision-making interface in practice? The answer is that we can combine the advantages of both interfaces. We can start by allowing users to put forward their hypothesis and evidence first to avoid automation bias [12]. Following this, they will be shown a ranking of hypotheses based on the level of uncertainty. This allows users to be aware of the AI’s recommendation (i.e., the hypothesis with the least uncertainty), as well as the supporting and opposing evidence for all possible hypotheses. Alternatively, we can use conformal prediction [75] to present multiple hypotheses within a given confidence bound. Uncertainty information can also reduce the number of possible hypotheses when there are too many to choose from, helping users focus on the most likely ones and avoid decision fatigue.
Moreover, the decision-aid can allow argumentation between the decision-maker and the AI. Users can provide their thoughts, including the evidence supporting their decisions. The decision-support system will then compare the user’s evidence with the AI’s evidence, as well as the user’s hypothesis with the AI’s hypothesis. Users can also modify the AI’s evidence based on their domain knowledge. This comparison helps users evaluate their decisions against those of the AI model and contribute to improving the model when they believe it is incorrect. Some recent works have used argumentation-based approaches to better design AI decision support. For example, Chiang et al. [17] introduced devil’s advocate to challenge the AI recommendation or the majority opinion within a group. Alternatively, the devil’s advocate can be used to present counter-arguments against the user’s decision [58]Ma et al. [57] proposed the idea of Deliberative AI, which allows the human user and the AI to deliberate conflicting evidence and arguments.
Regarding the presentation of evidence, it is not only about how much a feature contributes to the hypothesis, but also about whether the feature is important or relevant to it. In particular, even if we have strong evidence supporting a hypothesis, if the evidence is not relevant, its weight should be calibrated or even ignored when making the final decision.
6.4.2 Human Experiment Design
In both domains, we should control the cognitive load across experimental conditions by including a cognitive load measurement such as NASA-TLX scales [27] or reaction-time analysis. Since the hypothesis-driven condition can require more cognitive effort than the traditional recommendation-driven condition, it is important to determine whether any performance differences are due to cognitive load or the nature of the approach itself. Future work can consider using these cognitive load measures when evaluating different decision-support approaches. Another aspect to improve the human experiments is assessing whether trust calibration and decision quality improve over time. Current experiments only compare the decision-support approaches in short-term settings, without considering long-term effects of repeated exposure.
7 Conclusion
In this paper, we introduce the implementation of a hypothesis-driven approach by extending the Weight of Evidence framework. The new approach is evaluated in two domains: housing price prediction and skin cancer diagnosis. We found that the hypothesis-driven approach can help participants make better decisions by providing a more complete picture of the underlying decision aid. However, the recommendation-driven approach is more streamlined and easier to use, especially for beginners. Different decision-support approaches have their own pros and cons, but can be combined to provide a better decision-support tool in the future.
Acknowledgements
This research was supported by (a) the University of Melbourne Research Scholarship (MRS), (b) Australian Research Council Discovery Grant DP190103414 and undertaken using the (c) LIEF HPC-GPGPU Facility hosted at the University of Melbourne. This Facility was established with the assistance of LIEF Grant LE170100200.
Declaration of Competing Interest
H. Peter Soyer is a shareholder of MoleMap NZ Limited and e-derm consult GmbH and undertakes regular teledermatological reporting for both companies. H. Peter Soyer is a Medical Consultant for Canfield Scientific Inc and a Medical Advisor for First Derm.
Appendix A Appendix
Supplementary materials are available online.